OpenStudy (chris215):

Find f[g(x)] if f(x) = x^2 + 1 and g(x) = x^5

OpenStudy (dayakar):

f(x)= x^2 ---------(1) g(x) = x^5 ---------(2) f[g(x)] = f[ x^5] [from (2)]

OpenStudy (dayakar):

do u understand this

OpenStudy (chris215):

yes

OpenStudy (dayakar):

can u find remaining part

OpenStudy (chris215):

do I get (x^5+1)^2 ?

OpenStudy (dayakar):

r u doing

OpenStudy (dayakar):

f(x) = x^2 + 1-------(1) f( x^5) =( x^5) ^2 +1 [replacing x with x^5 in (1) ] = x^10 +1

OpenStudy (chris215):

ohh okk I get it now thank you!